xneg3 y2 zneg4 what is the exponent?
Answer by nyc_function(2741)

x^(-3)*y^2*z^(-4) =

x^(-3) can be written as 1/x^3

AND
z^(-4) can be written as 1/z^4

(1/x^3)(y^2)(1/z^4) = y^2/(x^3)(z^4)